Hello,
It costs 55c for domestic mail within Australia and $2.05 to post a letter to Europe. If I spent $10.00 to post 10 letters, how many of each type did I post?
Could you assist by providing algebra equation?
Thank you,
Bridgit

Answer by rothauserc(4718)

let x be the number of .55 cent letters and y be the number of $2.05 letters, then
x +y = 10 and
.55x +(2.05)y = 10.00
solve first equation for x and we get x = 10 -y, then substitute in second equation
.55(10-y) +(2.05)y = 10.00
5.50 -.55y +(2.05)y = 10.00
(1.5)y = 4.50
y = 3 and x = 7
we have 7 letters at .55 cents each and 3 letters at $2.05 each
let's check the answer
7(.55) + 3(2.05) = 10.00
3.85 + 6.15 = 10.00
10.00 = 10.00
our answer checks :-)